Find the selling price per pound of a tea mixture made from 40lb of tea costing $5.40 per pound and 60lb of tea costing $3.25 per pound.
Answer by checkley75(3666)

40*5.40+60*3.25=100X
216+195=100X
411=100X
X=411/100
X=$4.11 PER POUND OF THE MIX.